A superconducting material has a critical temperature of 4.1 K at zero magnetic

field. The critical field at 0 K is 0.04 Tesla. Determine the critical field at 2 K.


1
Expert's answer
2021-02-22T10:27:57-0500
"H_c=H_0\left(1-\left(\frac{T}{T_c}\right)^2\right)\\
H_c=0.04\left(1-\left(\frac{2}{4.1}\right)^2\right)=0.03\ T"